OpenStudy (dancergirl45):

Which of the following represents the factored form of f(x) = x3 − 81x? f(x) = x(x − 9)2 f(x) = (x + 9)(x − 9) f(x) = x(x + 9)(x − 9) f(x) = x(x2 − 9)

OpenStudy (dancergirl45):

@SerenityJustice74

OpenStudy (aveline):

Any idea?

OpenStudy (dancergirl45):

umm B?

OpenStudy (supercalifragalisticexspeaalli):

do u still need help @@dancergirl45

OpenStudy (dancergirl45):

yes

OpenStudy (supercalifragalisticexspeaalli):

@dancergirl45

OpenStudy (aveline):

Sorry lost internet connection. First you need to factor out an x =(x)(x^2-81) Now you simplify x^2-81 =(x)(x+9)(x-9)

OpenStudy (aveline):

B is just x^2-81, so it can't be B

OpenStudy (supercalifragalisticexspeaalli):

it can't be d either so the answer has to beee

OpenStudy (supercalifragalisticexspeaalli):

@dancergirl45

OpenStudy (dancergirl45):

C

OpenStudy (supercalifragalisticexspeaalli):

yepppp

OpenStudy (dancergirl45):

Thank you guys

OpenStudy (supercalifragalisticexspeaalli):

np